OpenStudy (anonymous):

You are given a box of 100 silver dollars, all facing heads up. You are instructed to shake the box 2 times; after each shake, you will remove all the dollars that are heads up before shaking again. You may keep all the dollars that are still tails up following the sixth shake. How many dollars will you most likely get to keep? 6 12 25 50

OpenStudy (anonymous):

probability of a dollar remaining a heads = 1/2 Expected no. of tails remaining therefore after 1st shake = 1/2 * 100 = 50 similaryl after second shake 25

OpenStudy (anonymous):

Thanks